Integrate show all work please ( no steps missing). 10) S cotx CSC3 Х dx
Answer #1

I=S cota dx =S cos(x) x sin 3 x da csc3x sin(a) I - S Cos (x) sin? (xx) dx het y = sin (x) dy= cos(x) dx Is y3 dy 4 flcc c is